Express the confidence interval

0.444

less thanpless than0.888in the form

ModifyingAbove p with caret plus or minus Upper E

.
Answer #1

Solution:

Given:

( 0.444 < p < 0.888 )

Lower Bound = 0.444

Upper Bound = 0.888

We have to express this confidence interval in the form of \hat{p} \pm E

\hat{p} = \frac{Lower \: \: Bound + Upper\: \: Bound }{2}

\hat{p} = \frac{ 0.444 + 0.888 }{2}

\hat{p} = \frac{ 1.332 }{2}

\hat{p} =0.666

and

E = \frac{Upper \: \: Bound -Lower \: \: Bound }{2}

E = \frac{0.888 - 0.444}{2}

E = \frac{0.444 }{2}

E = 0.222

Thus we get:

\hat{p} \pm E

\mathbf{{\color{DarkOrange} 0.666 \pm 0.222}}